# Fast Policy Learning for Linear-Quadratic Control with Entropy Regularization

### v3

@@ 6 Application: Transfer Learning
$\left|C^{\overline{\mathcal{M}}}(K^{*},\Sigma^{*})-C^{{\mathcal{M}}}(K^{*},\Sigma^{*})\right|$ (40)
$\leq$ $2c_{\gamma,\rho}\Big{(}\|D_{0}+W\|_{q}+\frac{\gamma}{1-\gamma}+1\Big{)}\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}\Big{(}\|A-\overline{A}\|_{q}+\|B-\overline{B}\|_{q}\Big{)}.$
Finally, plugging (39)-(40) into (35), we have
$\left|C^{\overline{\mathcal{M}}}(\overline{K}^{*},\overline{\Sigma}^{*})-C^{\overline{\mathcal{M}}}(K^{*},\Sigma^{*})\right|$
$\leq$ $4c_{\gamma,\rho}\Big{(}\|D_{0}+W\|_{q}+\frac{\gamma}{1-\gamma}+1\Big{)}\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}\Big{(}\|A-\overline{A}\|_{q}+\|B-\overline{B}\|_{q}\Big{)}.$
Theorem 6.1 suggests that if the environments $\mathcal{M}$ and $\overline{\mathcal{M}}$ are sufficiently close in the sense of (34), then $(K^{*},\Sigma^{*})$ serves an efficient initial policy for $\overline{\mathcal{M}}$ which directly leads to a super-linear convergence for the new learning problem.
@@ 7 Numerical Experiments
This section provides numerical experiments using (RPG) and (IPO) to illustrate the theoretical results established in Section 4, 5, and 6.
Setup.
@@
(1) Parameters: $n=400$, $k=200$, $A$, $B$, $Q$, $R$ are generated randomly. The scaling of $A$ is chosen so that $A$ is stabilizing with high probability ($\sigma_{\max}(A)<\frac{1}{\sqrt{\gamma}}$). $Q$ and $R$ are generated to be positive definite. The regularization parameter $\tau$ is chosen to be $\sigma_{\min}(R)$. Initialization: $K^{(0)}_{i,j}=0.01$ for all $i,j$, $\Sigma^{(0)}=I.$
@@
(2) Transfer learning setup: $\overline{A}$ and $\overline{B}$ are the state transition matrices which are generated by adding a perturbation to $A$ and $B$: $\overline{A}_{i,j}=A_{i,j}+u_{i,j}$, $\overline{B}_{i,j}=B_{i,j}+u^{\prime}_{i,j}$, where $u_{i,j}$ and $u^{\prime}_{i,j}$ are sampled from a uniform distribution on $[0,10^{-3}]$. The initialization of $\overline{K}$ and $\overline{\Sigma}$ are the optimal solution $K^{*}$ and $\Sigma^{*}$ with state transition matrices $A$ and $B$.
Figure 1: Performances of (0(a)) (RPG); (0(b)) (IPO); (0(c)) transfer learning using (IPO) with $(\overline{K}^{(0)},\overline{\Sigma}^{(0)})=(K^{*},\Sigma^{*})$ and state transitions $(\overline{A},\overline{B})$.
Performance measure.
@@
We use the following normalized error to quantify the performance of a given policy ${K,\Sigma}$,
[…]
where ${K}^{*},\Sigma^{*}$ is the optimal policy defined in Theorem 2.1.
(Fast) Convergence.
@@
Figure 0(a) shows the linear convergence of (RPG), and Figure 0(b) shows the superior convergence rate of (IPO). The normalized error falls below $10^{-14}$ within just 6 iterations, and from the third iteration, it enters a region of super-linear convergence. Figure 0(c) shows the result of applying transfer learning using (IPO) in a perturbed environment, when the optimal policy in Figure 0(a) and 0(b) serves as an initialization. Figure 0(c) shows that if the process commences within a super-linear convergence region, then the error falls below $10^{-12}$ in just 2 epochs.
@@ 8.1 Proofs in Section 2
Proof.
[…]
where $\mathcal{L}(u,v,\lambda):=(u^{\top}Mu+b^{\top}u)v+\tau v\log v+\lambda v.$ Since $\frac{\partial\mathcal{L}}{\partial{v}}=\lambda+u^{\top}Mu+b^{\top}u+\tau+\tau\log v,$ by setting $\frac{\partial\mathcal{L}}{\partial{v}}=0,$ we then get $v\propto e^{-\frac{1}{\tau}(u^{\top}Mu+b^{\top}u)}.$ Thus, the optimal solution $p^{*}$ is a multivariate Gaussian distribution with $\mathcal{N}\left(-\frac{1}{2}M^{-1}b,\frac{\tau}{2}M^{-1}\right)$.
@@ 8.2 Proofs in Section 3
$\nabla_{\Sigma}C(K,\Sigma)$ in (18) can be checked by direct gradient calculation. To verify $\nabla_{K}C(K,\Sigma)$ in (18), first define $f:\mathcal{X}\times\mathbb{R}^{k\times n}\rightarrow\mathbb{R}$ by $f(y,K):=y^{\top}P_{K}y,\forall y\in\mathcal{X}$ and we aim to find the gradient of $f$ with respect to $K$. For any $y\in\mathcal{X},$ by the Riccati Equation for $P_{K}$ in (17) we have
@@
$\nabla_{K}f((A-BK)y,K)$ has two terms: one with respect to the input $(A-BK)y$ and one with respect to $K$ in the subscript of $P_{K}$. This implies
$\nabla_{K}f(y,K)=2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)yy^{\top}+\gamma\nabla_{K}f(y^{\prime},K)|_{y^{\prime}=(A-BK)y}$ (41)
@@
Since $C(K,\Sigma)=\mathbb{E}_{x_{0}\sim\mathcal{D}}[x_{0}P_{K}x_{0}]+q_{K,\Sigma}$ with $P_{K}$ and $q_{K,\Sigma}$ satisfying (17), then the gradient of $C(K,\Sigma)$ with respect to $K$ is
@@
=$\overset{(a)}{=}$ $\mathbb{E}\Big{[}2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)x_{0}x_{0}^{\top}+\gamma\nabla_{K}f(\hat{x}_{1},K)|_{\hat{x}_{1}=(A-BK)x_{0}}$
@@
=$\overset{(b)}{=}$ $\mathbb{E}\Big{[}2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)x_{0}x_{0}^{\top}+\gamma\nabla_{K}f({x}_{1},K)|_{{x}_{1}=A+Bu_{0}+w_{0}}$
@@
=$\overset{(c)}{=}$ $2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)\mathbb{E}\left[\sum_{i=0}^{\infty}\gamma^{i}x_{i}x_{i}^{\top}\right],$
where $(a)$ follows from applying (41) with $y=x_{0}$ and taking the gradient of $q_{K,\Sigma}$ in (17) with respect to $K$; $(b)$ follows from
$\mathbb{E}_{x_{0},u_{0},w_{0}}[f(Ax_{0}+Bu_{0}+w_{0}),K]=\mathbb{E}_{x_{0},w_{0}}[f\left((A-BK)x_{0},K\right)+w_{0}^{\top}P_{K}w_{0}]+\Sigma B^{\top}P_{K}B.$
Using recursion to get $(c)$. ∎
@@
For a given policy $\pi_{K,\Sigma}(\cdot|x)=\mathcal{N}(-Kx,\Sigma)$ with parameter $K$ and $\Sigma$, we define the state-action value function (also known as $Q$-function) $Q_{K,\Sigma}:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$ as the cost of the policy starting with $x_{0}=x$, taking a fixed action $u_{0}=u$ and then proceeding with $\pi_{K,\Sigma}$. The $Q$-function is related to the value function $J_{K,\Sigma}$ defined in (2) as
$Q_{K,\Sigma}(x,u)=$ $x^{\top}Qx+u^{\top}Ru+\tau\pi_{K,\Sigma}(u|x)+\gamma\mathbb{E}\left[J_{K,\Sigma}(Ax+Bu+w)\right],$ (42)
for any $(x,u)\in\mathcal{X}\times\mathcal{A}$. By definition of the $Q$-function, we also have the relationship:
$\forall x\in\mathcal{X}:\quad$ $J_{K,\Sigma}(x)=\mathbb{E}_{u\sim\pi(\cdot|x)}\left[Q_{K,\Sigma}(x,u)\right].$ (43)
We also introduce the advantage function $A_{K,\Sigma}:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$ of the policy $\pi$:
$A_{K,\Sigma}(x,u)=Q_{K,\Sigma}(x,u)-J_{K,\Sigma}(x),$ (44)
@@
With the notations of the Q-function in (42) and the advantage function in (44), we first provide a convenient form for the difference of the cost functions with respect to two different policies in Lemma 8.1. This will be used in the proof of Lemma 3.2.
@@
Suppose policies $\pi$ and $\pi^{\prime}$ are in form of (14) with parameters $(K^{\prime},\Sigma^{\prime})$ and $(K,\Sigma).$ Let $\{x_{t}^{\prime}\}_{t=0}^{\infty}$ and $\{u_{t}^{\prime}\}_{t=0}^{\infty}$ be state and action sequences generated by $\pi^{\prime}$ with noise sequence $\{w_{t}^{\prime}\}_{t=0}^{\infty}$ (i.i.d with mean 00 and covariance $W$), i.e., $x_{t+1}^{\prime}=Ax_{t}^{\prime}+Bu_{t}^{\prime}+w_{t}^{\prime}$. Then for any $x\in\mathcal{X},$
$J_{K^{\prime},\Sigma^{\prime}}(x)-J_{K,\Sigma}(x)=\mathbb{E}_{\pi^{\prime}}\left[\left.\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x_{t}^{\prime},u_{t}^{\prime})\right|x_{0}^{\prime}=x\right],$ (45)
@@
$\mathbb{E}_{\pi^{\prime}}\left[A_{K,\Sigma}(x,u)\right]=x^{\top}(K^{\prime}-K)^{\top}(R+\gamma B^{\top}P_{K}B)(K^{\prime}-K)x$ (46)
@@
=$\overset{(a)}{=}$ $\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(c_{t}^{\prime}-J_{K,\Sigma}(x_{t}^{\prime})\right)+\sum_{t=1}^{\infty}\gamma^{t}J_{K,\Sigma}(x_{t}^{\prime})\right]=\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(c_{t}^{\prime}+\gamma J_{K,\Sigma}(x_{t+1}^{\prime})-J_{K,\Sigma}(x_{t}^{\prime})\right)\right]$
@@
where $(a)$ follows from $x_{0}^{\prime}=x$. By (42) and (44), we can continue the above calculations by
@@
Plug (17) into the above equation, then use (16) and write $K^{\prime}=K+K^{\prime}-K$ to get
@@
Plugging (17) finishes the proof to verify (46). ∎
Proof.
(of Lemma 3.2). By Lemma 8.1, for any $\pi$ and $\pi^{\prime}$ in form of (14) with parameter $(K,\Sigma)$ and $(K^{\prime},\Sigma^{\prime})$ respectively, and for any $x\in\mathcal{X}$,

### v4

@@ 6 Transfer Learning for RL
(6.7) $\left|\overline{C}(K^{*},\Sigma^{*})-C(K^{*},\Sigma^{*})\right|$
$\leq 2c_{\gamma,\rho}\Big(\|D_{0}+W\|_{q}+\frac{\gamma}{1-\gamma}+1\Big)\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}\Big(\|A-\overline{A}\|_{q}+\|B-\overline{B}\|_{q}\Big).$
Finally, plugging (6.6)-(6.7) into (6.3), we have
(6.8) $\left|\overline{C}(\overline{K}^{*},\overline{\Sigma}^{*})-\overline{C}(K^{*},\Sigma^{*})\right|$
$\leq 4c_{\gamma,\rho}\Big(\|D_{0}+W\|_{q}+\frac{\gamma}{1-\gamma}+1\Big)\frac{\|Q\|+\|R\|}{1-\gamma\rho^{2}}\Big(\|A-\overline{A}\|_{q}+\|B-\overline{B}\|_{q}\Big).$
7 Model-free Extension
Model-based convergence provides a foundation for model-free analysis, as demonstrated [15, 21] where the more challenging policy convergence analysis for the model-based setting is followed by a relatively routine sample-based analysis of zeroth-order gradient approximation [35, 30, 3, 14]. Similarly, our analysis of local superlinear convergence and transfer learning applications within a model-based framework can be extended to developing a model-free algorithm, for instance, Algorithm 1.
Algorithm 1 Policy Gradient Estimation with Unknown Parameters
Input: $K,\Sigma$, the number of trajectories $m$, smoothing parameter $r$, dimension $D_{K}$ and $D_{\Sigma}$
 Apply Cholesky decomposition to matrix $\Sigma$ to get $L$ such that $\Sigma=LL^{\top}$.
for $i=1,2,\cdots,m$ do
  Sample a policy $\widehat{\Sigma}_{i}=\widehat{L}_{i}(\widehat{L}_{i})^{\top}$, where ${\text{vec}}(\widehat{L}_{i})={\text{vec}}(L)+U_{i}$, where $U_{i}$ is drawn uniformly at random over vectors in $\mathbb{R}^{D_{\Sigma}}$ such that $\|U_{i}\|_{F}=r.$ Simulate with policy $(K,\widehat{\Sigma}_{i})$ from $x_{0}\sim\mathcal{D}$ for $l$ steps. Let $\widehat{C}_{i}$ denote the empirical estimates:
$\widehat{C}_{i}=\sum_{t=1}^{l}\gamma^{t}c_{t},$
where $c_{t}$ and $x_{t}$ are the costs and states on this trajectory.
  Sample a policy $\widehat{K}_{i}=K+U_{i}^{\prime}$, where $U_{i}^{\prime}$ is drawn uniformly at random over matrices in $\mathbb{R}^{k\times n}$ such that $\|U_{i}^{\prime}\|_{F}=r$. Simulate with policy $(\widehat{K}_{i},{\Sigma}_{i})$ from $x_{0}\sim\mathcal{D}$ for $l$ steps and let $\widehat{C_{i}^{\prime}}$ and $\widehat{S_{i}^{\prime}}$ denote the empirical estimates:
$\widehat{C_{i}^{\prime}}=\sum_{t=1}^{l}\gamma^{t}c_{t}^{\prime},\quad\widehat{S_{i}^{\prime}}=\sum_{t=1}^{l}\gamma^{t}x_{t}^{\prime}x_{t}^{\top}.$
end for
return the estimates of $\nabla_{K}C(K,\Sigma),\nabla_{\Sigma}C(K,\Sigma),S_{K,\Sigma}$:
$\widehat{\nabla_{{\text{vec}}(\Sigma)}C(K,\Sigma)}$ $=\left(\nabla_{{\text{vec}}(L)}{\text{vec}}(\widehat{\Sigma}(\widehat{L}))^{\top}\right)^{-1}\cdot\left(\frac{1}{m}\sum_{i=1}^{m}\frac{D_{\Sigma}}{r^{2}}\widehat{C}_{i}U_{i}\right),$
$\widehat{\nabla_{K}C(K,\Sigma)}$ $=\frac{1}{m}\sum_{m=1}^{l}\gamma^{t}\frac{D_{K}}{r^{2}}\widehat{C_{i}^{\prime}}U_{i}^{\prime},\quad\widehat{S_{K,\Sigma}}=\frac{1}{m}\sum_{i=1}^{m}\widehat{S_{i}^{\prime}}.$
In the setting with unknown model parameters $A,B,Q,R$, where the controller has only simulation access to the model, we apply a zeroth-order optimization method to approximate the gradient $\widehat{\nabla_{K}C(K,\Sigma)}$ and $\widehat{\nabla_{\Sigma}C(K,\Sigma)}$, as in Algorithm 1.
The updating rule for $\widehat{\nabla_{K}C(K,\Sigma)}$ is standard [15, 21]. We now explain the expression for $\widehat{\nabla_{\Sigma}C(K,\Sigma)}$. By Lemma 4.5 and Theorem 4.2, $\Sigma$ is positive definite (with the time index $t$ omitted for ease of exposition). Let $L$ denote the Cholesky decomposition of $\Sigma$ such that $\Sigma=LL^{\top}.$ Let ${\text{vec}}(\Sigma)\in\mathbb{R}^{D_{\Sigma}}$ and ${\text{vec}}(L)\in\mathbb{R}^{D_{\Sigma}}$ denote the stacked vectors of the lower-triangular entries in matrices $\Sigma\in\mathbb{R}^{k\times k}$ and $L\in\mathbb{R}^{k\times k}$ respectively, with $D_{\Sigma}:=\frac{k(k+1)}{2}$. Similarly, denote ${\text{vec}}(K)\in\mathbb{R}^{D_{K}}$ as the stacked vectors of all entries in $K\in\mathbb{R}^{k\times n}$, with $D_{K}:=k\times n.$ Then $\nabla_{{\text{vec}}(L)}C(K,\Sigma)=\nabla_{{\text{vec}}(L)}{\text{vec}}(\Sigma(L))^{\top}\cdot\nabla_{{\text{vec}}(\Sigma)}C(K,\Sigma)$. In Algorithm 1, we approximate $\nabla_{{\text{vec}}(L)}C(K,\Sigma)$ with zeroth-order estimate and the above equation to get $\widehat{\nabla_{{\text{vec}}(\Sigma)}C(K,\Sigma)}$. The estimate $\widehat{\nabla_{\Sigma}C(K,\Sigma)}$ can be obtained by rearranging the entries of $\widehat{\nabla_{{\text{vec}}(\Sigma)}C(K,\Sigma)}$ into a matrix form.
@@ 8 Numerical Experiments
This section provides numerical experiments using (RPG) and (IPO) to illustrate the results established in Section 4, 5, and 6.
Setup
@@
(1) Parameters: $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times k}$, $Q\in\mathbb{S}^{n}_{+}$, $R\in\mathbb{S}^{k}_{++}$ are generated randomly. The scaling of $A$ is chosen so that $A$ is stabilizing with high probability ($\sigma_{\max}(A)<\frac{1}{\sqrt{\gamma}}$). Initialization: $K^{(0)}_{i,j}=0.01$ for all $i,j$, $\Sigma^{(0)}=I.$
@@
(2) Transfer learning setup: $\overline{A}$ and $\overline{B}$ are the state transition matrices which are generated by adding a perturbation to $A$ and $B$: $\overline{A}_{i,j}=A_{i,j}+u_{i,j}$, $\overline{B}_{i,j}=B_{i,j}+u^{\prime}_{i,j}$, where $u_{i,j}$ and $u^{\prime}_{i,j}$ are sampled from a uniform distribution on $0,10^{-3}]$. The initialization of $\overline{K}$ and $\overline{\Sigma}$ are the optimal solution $K^{*}$ and $\Sigma^{*}$ with state transition matrices $A$ and $B$.
Performance measure
@@
We use normalized error to quantify the performance of a given policy $K,\Sigma$}, i.e., $\textsc{normalized error }=\frac{C({K,\Sigma})-C\left({K}^{*},\Sigma^{*}\right)}{C\left({K}^{*},\Sigma^{*}\right)},$ where ${K}^{*},\Sigma^{*}$ is the optimal policy defined in Theorem 2.1.
(Fast) Convergence
@@
Figure 1(a) shows the linear convergence of (RPG), and Figure 1(b) shows the superior convergence rate of (IPO). The normalized error falls below $10^{-14}$ within just 6 iterations, and from the third iteration, it enters a region of super-linear convergence. Figure 1(c) shows the result of applying transfer learning using (IPO) in a perturbed environment, when the optimal policy in Figure 1(a) and 1(b) serves as an initialization. Figure 1(c) shows that if the process commences within a super-linear convergence region, then the error falls below $10^{-12}$ in just two epochs.
Figure 1: Performances of (1(a)) (RPG); (1(b)) (IPO); (1(c)) transfer learning using (IPO) with $(\overline{K}^{(0)},\overline{\Sigma}^{(0)})=(K^{*},\Sigma^{*})$ and state transitions $(\overline{A},\overline{B})$. $n=400$, $k=200.$ The regularization parameter $\tau$ is chosen to be $\sigma_{\min}(R)$.
Figure 2: Regularized Policy Gradient (RPG) with different regularization parameters $\tau$. Left: $n=200,k=10;$ Right: $n=200,k=50.$
Figure 3: Comparison between Iterative Policy Optimization (IPO) and Gauss-Newton update on $K$ with constant covariance matrix (8.1). $n=200,k=50$. Left: $\tau=0.01;$ Right: $\tau=0.5.$
Regularization parameter $\tau$
To demonstrate that entropy regularization can accelerate convergence, we conduct experiments with (RPG) under two settings: $n=200,k=10$ and $n=200,k=50$. We run (RPG) using various values of $\tau$. Figure 2 illustrates that, in both settings, a larger value of $\tau$ results in a faster linear convergence rate to the optimal solution of (3.3). These results confirm that increasing the regularization parameter enhances the convergence speed, highlighting the practical benefits of entropy regularization in achieving faster optimization.
The importance of updating $\Sigma$
As discussed in Section 5, updating $K$ in (IPO) is identical to updating step of the Gauss-Newton algorithm in an unregularized setting. However, unlike the Gauss-Newton algorithm, (IPO) also updates the covariance matrix $\Sigma$ simultaneously. To see the effect of dynamic updating of $\Sigma$, we compare the performance of (IPO) with the Gauss-Newton update on $K$ with a constant covariance matrix $\Sigma$ given by:
(8.1) $K^{(t+1)}$ $=K^{(t)}-\left(R+\gamma B^{\top}P_{K^{(t)}}B\right)^{-1}E_{K^{(t)}},$
$\Sigma^{(t+1)}$ $=\sigma I,$
where $\sigma$ is a fixed positive scalar. Figure 3 illustrates that the (IPO) algorithm with updates of $\Sigma$ achieves a noticeably faster and superlinear convergence rate, compared to the Gauss-Newton update with a fixed $\Sigma$. This dynamic update of $\Sigma$ allows (IPO) to reach the optimal point of Problem (3.3) more efficiently, highlighting the importance of adapting the covariance matrix during iterations.
@@ 9.1 Proofs in Section 2
9.1.1 Proof of Lemma 2.2
Denote the domain of the decision variable as $\mathcal{X}=\{p:\mathcal{A}\mapsto[0,\infty)\},$ and the feasible set as $\mathcal{F}=\{p:\mathcal{A}\mapsto[0,\infty)\mid\int_{\mathcal{A}}p(u)\mathrm{d}u=1\}\subseteq\mathcal{X}.$ Let $f:\mathcal{X}\mapsto\mathbb{R}$ denote the objective function, i.e.,
$f(p)=\mathbb{E}_{u\sim p(\cdot)}\left[u^{\top}Mu+b^{\top}u+\tau\log p(u)\right].$
Let $\lambda$ be a Lagrangian multiplier to the constraint $\int_{\mathcal{A}}p(u)\mathrm{d}u=1.$ Consider
$\mathcal{L}(p,\lambda)$ $=\int_{\mathcal{A}}\left(u^{\top}Mu+b^{\top}u+\tau\log p(u)\right)p(u)\mathrm{d}u+\lambda\left(\int_{\mathcal{A}}p(u)\mathrm{d}u-1\right)$
$=\int_{\mathcal{A}}{L}\left(u,p(u),\lambda\right)\mathrm{d}u-\lambda,$
where ${L}(u,v,\lambda):=(u^{\top}Mu+b^{\top}u)v+\tau v\log v+\lambda v.$ Additionally, define $g(\lambda)=\inf_{\mathcal{X}}\mathcal{L}(p,\lambda)$.
We now show the strong duality result:
(9.1) $g(\lambda^{*})=\inf_{p\in\mathcal{F}}f(p),$
with $\lambda^{*}=\arg\max_{\lambda\in\mathbb{R}}g(\lambda).$
First, the weak duality result follows from
(9.2) $g(\lambda)=\inf_{p\in\mathcal{X}}\mathcal{L}(p,\lambda)\leq\inf_{p\in\mathcal{F}}\mathcal{L}(p,\lambda)=\inf_{p\in\mathcal{F}}f(p),\text{ for any }\lambda\in\mathbb{R}.$
Moreover, since $\frac{\partial{L}}{\partial{v}}(u,v,\lambda)=\lambda+u^{\top}Mu+b^{\top}u+\tau+\tau\log v,$ for any $\lambda\in\mathbb{R},u\in\mathcal{A}$, the minimizer $p_{\lambda}(u)$ of $L(u,\cdot,\lambda)$ satisfies
(9.3) $p_{\lambda}(u)=\exp\left(-\frac{1}{\tau}(\lambda+u^{\top}Mu+b^{\top}u)-1\right).$
Therefore, by applying (9.3) to (9.2), we have
(9.4) $g(\lambda)=\mathcal{L}(p_{\lambda},\lambda)=-\tau\left(\exp\left(-\frac{\lambda}{\tau}-1\right)\cdot C+\frac{\lambda}{\tau}\right),$
where $C:=\int_{\mathcal{A}}\exp\left(-\frac{1}{\tau}(u^{\top}Mu+b^{\top}u)\right)\mathrm{d}u$. Direct computation yields the maximizer of $g$ in (9.2) as $\lambda^{*}=\tau\log C-\tau$. Plugging $\lambda^{*}$ to (9.3) shows $\int_{\mathcal{A}}p_{\lambda^{*}}(u)\mathrm{d}u=1,$ implying $p_{\lambda^{*}}\in\mathcal{F}$ and the strong duality (9.1) holds. Finally, by (9.1) and (9.3), it is clear that the optimal solution is a multivariate Gaussian distribution with $\mathcal{N}\left(-\frac{1}{2}M^{-1}b,\frac{\tau}{2}M^{-1}\right)$.
@@ 9.2 Proofs in Section 3
$\nabla_{\Sigma}C(K,\Sigma)$ in (3.6) can be checked by direct gradient calculation. To verify $\nabla_{K}C(K,\Sigma)$ in (3.6), first define $f:\mathcal{X}\times\mathbb{R}^{k\times n}\rightarrow\mathbb{R}$ by $f(y,K):=y^{\top}P_{K}y,\forall y\in\mathcal{X}$ and we aim to find the gradient of $f$ with respect to $K$. For any $y\in\mathcal{X},$ by the Riccati equation for $P_{K}$ in (3.5) we have
@@
$\nabla_{K}f((A-BK)y,K)$ has two terms: one with respect to the input (ABK)y(A-BK)y and one with respect to $K$ in the subscript of $P_{K}$. This implies
(9.5) $\nabla_{K}f(y,K)=2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)yy^{\top}+\gamma\nabla_{K}f(y^{\prime},K)|_{y^{\prime}=(A-BK)y}$
@@
Since $C(K,\Sigma)=\mathbb{E}_{x_{0}\sim\mathcal{D}}[x_{0}P_{K}x_{0}]+q_{K,\Sigma}$ with $P_{K}$ and $q_{K,\Sigma}$ satisfying (3.5), then the gradient of $C(K,\Sigma)$ with respect to $K$ is
@@
$\overset{(a)}{=}$ $\mathbb{E}\Big[2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)x_{0}x_{0}^{\top}+\gamma\nabla_{K}f(\hat{x}_{1},K)|_{\hat{x}_{1}=(A-BK)x_{0}}$
@@
$\overset{(b)}{=}$ $\mathbb{E}\Big[2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)x_{0}x_{0}^{\top}+\gamma\nabla_{K}f({x}_{1},K)|_{{x}_{1}=A+Bu_{0}+w_{0}}$
@@
$\overset{(c)}{=}$ $2\left(-\gamma B^{\top}P_{K}(A-BK)+RK\right)\mathbb{E}\left[\sum_{i=0}^{\infty}\gamma^{i}x_{i}x_{i}^{\top}\right],$
where $(a)$ follows from applying (9.5) with $y=x_{0}$ and taking the gradient of $q_{K,\Sigma}$ in (3.5) with respect to $K$; $(b)$ follows from
$\mathbb{E}_{x_{0},u_{0},w_{0}}[f(Ax_{0}+Bu_{0}+w_{0}),K]=\mathbb{E}_{x_{0},w_{0}}[f\left((A-BK)x_{0},K\right)+w_{0}^{\top}P_{K}w_{0}]+\Sigma B^{\top}P_{K}B.\vskip-4.2679pt$
Using recursion to get $(c)$.
@@
For a given policy $\pi_{K,\Sigma}(\cdot|x)=\mathcal{N}(-Kx,\Sigma)$ with parameter $K$ and $\Sigma$, we define the state-action value function (also known as $Q$-function) $Q_{K,\Sigma}:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$ as the cost of the policy starting with $x_{0}=x$, taking a fixed action $u_{0}=u$ and then proceeding with $\pi_{K,\Sigma}$. The $Q$-function is related to the value function $J_{K,\Sigma}$ defined in (2.1) as
(9.6) $Q_{K,\Sigma}(x,u)=$ $x^{\top}Qx+u^{\top}Ru+\tau\pi_{K,\Sigma}(u|x)+\gamma\mathbb{E}\left[J_{K,\Sigma}(Ax+Bu+w)\right],$
for any $(x,u)\in\mathcal{X}\times\mathcal{A}$. By definition of the $Q$-function, we also have the relationship $J_{K,\Sigma}(x)=\mathbb{E}_{u\sim\pi(\cdot|x)}\left[Q_{K,\Sigma}(x,u)\right],\,\forall x\in\mathcal{X}.$ We then introduce the advantage function $A_{K,\Sigma}:\mathcal{X}\times\mathcal{A}\rightarrow\mathbb{R}$ of the policy $\pi$:
(9.7) $A_{K,\Sigma}(x,u)=Q_{K,\Sigma}(x,u)-J_{K,\Sigma}(x),\vskip-2.84526pt$
@@
With the notations of the Q-function in (9.6) and the advantage function in (9.7), we first provide a convenient form for the difference of the objective functions with respect to two different policies in Lemma 9.1. It takes a similar form as [15, Lemma 10], but with an additional term on $\Sigma$. This will be used in the proof of Lemma 3.2.
@@
Suppose policies $\pi$ and $\pi^{\prime}$ are in form of (3.2) with parameters $(K^{\prime},\Sigma^{\prime})$ and $(K,\Sigma).$ Let $\{x_{t}^{\prime}\}_{t=0}^{\infty}$ and $\{u_{t}^{\prime}\}_{t=0}^{\infty}$ be state and action sequences generated by $\pi^{\prime}$ with noise sequence $\{w_{t}^{\prime}\}_{t=0}^{\infty}$ (i.i.d with mean 0 and covariance $W$), i.e., $x_{t+1}^{\prime}=Ax_{t}^{\prime}+Bu_{t}^{\prime}+w_{t}^{\prime}$. Then for any $x\in\mathcal{X},$
(9.8) $J_{K^{\prime},\Sigma^{\prime}}(x)-J_{K,\Sigma}(x)=\mathbb{E}_{\pi^{\prime}}\left[\left.\sum_{t=0}^{\infty}\gamma^{t}A_{K,\Sigma}(x_{t}^{\prime},u_{t}^{\prime})\right|x_{0}^{\prime}=x\right],\vskip-2.84526pt$
@@
(9.9) $\mathbb{E}_{\pi^{\prime}}\left[A_{K,\Sigma}(x,u)\right]=x^{\top}(K^{\prime}-K)^{\top}(R+\gamma B^{\top}P_{K}B)(K^{\prime}-K)x$
@@
$\overset{(a)}{=}$ $\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(c_{t}^{\prime}-J_{K,\Sigma}(x_{t}^{\prime})\right)+\sum_{t=1}^{\infty}\gamma^{t}J_{K,\Sigma}(x_{t}^{\prime})\right]$
$=$ $\mathbb{E}_{\pi^{\prime}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(c_{t}^{\prime}+\gamma J_{K,\Sigma}(x_{t+1}^{\prime})-J_{K,\Sigma}(x_{t}^{\prime})\right)\right]$
@@
where $(a)$ follows from $x_{0}^{\prime}=x$. By (9.6) and (9.7), we can continue the above calculations by
@@
Plug (3.5) into the above equation, then use (3.4) and write $K^{\prime}=K+K^{\prime}-K$ to get
@@
Plugging (3.5) finishes the proof to verify (9.9).
Proof 9.3.
(of Lemma 3.2). By Lemma 9.1, for any $\pi$ and $\pi^{\prime}$ in form of (3.2) with parameter $(K,\Sigma)$ and $(K^{\prime},\Sigma^{\prime})$ respectively, and for any $x\in\mathcal{X}$,